# Perfect Divisibility and Coloring of Some Bull-Free Graphs

### v3

@@ Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edges, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [The Electron. J. of Combin. 29 (2022) P3.19.] proved that every (fork, bull)-free graph is perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307.] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. Let $H\in${{odd torch}, $\{P_{8},C_{5}\}\}$. In this paper, we prove that every ($H$, bull)-free graph is perfectly divisible. We also prove that a ($P_{6}$, bull)-free graph is perfectly divisible if and only if it contains no Mycielski-Grötzsch graph as an induced subgraph. As corollaries, these graphs are $\binom{\omega+1}{2}-$colorable. Notice that every odd torch contains an odd hole, an induced $P_{5}$, and an induced fork. Therefore, our results generalize their findings. Moreover, we prove that every ($P_{6}$, bull)-free graph $G$ satisfies $\chi(G)\leq\omega(G)^{7}$.
Key words and phrases: bull-free graphs; perfectly divisibility; chromatic number.
AMS 2000 Subject Classifications: 05C15, 05C75
@@
For a given positive integer $k$, we use the notation $[k]$ to denote the set $\{1,\ldots,k\}$. A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. non-adjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
[…]
An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole. A hole is odd if $k$ is odd, and even otherwise. An antihole is the complement graph of a hole. A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$. The famous Strong Perfect Graph Theorem [7] was established by Chudnovsky et al. in 2006:
@@
Let $F$ denote the Mycielski-Grötzsch graph $G_{4}$. A bull is a graph consisting of a triangle with two disjoint pendant edges, an odd torch is a graph obtained from an odd hole by adding an edge $xy$ such that $x$ is non-adjacent to any vertex on the odd hole and the set of neighbors of $y$ on the odd hole is a stable set, a fork is a graph obtained from $K_{1,3}$ by subdividing an edge once, and a E is a graph obtained from an induced $P_{5}=v_{1}v_{2}v_{3}v_{4}v_{5}$ by adding an new vertex exactly adjacent to $v_{3}$ in the $P_{5}$. (See Figure 1)
Figure 1: Illustration of some graphs.
A graph is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoáng in [12]. Chudnovsky and Sivaraman [9] proved that every (odd hole, bull)-free graph and every ($P_{5}$, bull)-free graph are perfectly divisible. Karthick et al. [15] proved that every (fork, bull)-free graph is perfectly divisible, and this class of graphs was generalized by Wu and Xu [19]. Chen and Xu [5] proved that every ($P_{7},C_{5}$, bull)-free graph is perfectly divisible. In this paper, we prove the following theorems.
@@
Let $G$ be a (oddtorch,bull)(odd~torch,bull)-free graph. Then $G$ is perfectly divisible.
@@
Let $G$ be a $(P_{8},C_{5},bull)$-free graph. Then $G$ is perfectly divisible.
It is easy to verify that every odd torch contains an odd hole, an induced $P_{5}$, an induced fork, and the graph $E$. Therefore, all the class of (odd hole, bull)-free graphs, ($P_{5}$, bull)-free graphs, (fork, bull)-free, and ($E$, bull)-free graphs are subclasses of the class of (odd torch, bull)-free graphs. Theorem 1.2 generalizes the results of Chudnovsky and Sivaraman [9] and Karthick et al. [15]; and Theorem 1.3 generalizes the result of Chen and Xu [5].
Recall that we use $F$ to denote the Mycielski-Grötzsch graph. It is certain that $F$ is ($P_{6}$, bull)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$.). In fact, the graph $G$ obtained from $F$ by adding a $K_{n}$ in which each vertex is adjacent to all the vertices in $F$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible. Since every ($P_{5}$, bull)-free graph is perfectly divisible, a natural problem is that under what conditions is ($P_{6}$, bull)-free graph perfectly divisible ? In this paper, we prove the following theorem.
@@
Let $G$ be a $(P_{6},\text{bull})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
@@
By a simple induction on $\omega(G)$, we have that $\chi(G)\leq\binom{\omega(G)+1}{2}$ for each perfectly divisible graph $G$. According to Theorems 1.2, 1.3 and 1.4, we can directly derive the following corollary.
@@
Let $G$ be an $(H,\text{bull})$-free graph, where $H\in\{\{odd~torch\},\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
Corollary 1.2
Let $G$ be an ((odd torch, $C_{3}$))-free graph. Then $\chi(G)\leq 3$.
We also consider the binding function of ($P_{6}$, bull)-free graphs, and prove the following theorem.
@@
Let $G$ be a $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
Let $G$ be a graph and let $w$ be a weight function on $V(G)$. We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$. A graph $G$ is perfectly weight divisible [9, 12] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$, $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$. It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible. If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short).
As mentioned above, the structure and coloring problem of bull-free graphs has garnered significant attention, with new methods and research content [4, 6, 11, 13, 18] being continuously proposed. This paper utilizes the concept of perfectly weight divisibility to further analyze the structure of bull-free graphs and bull-free MNWD graphs, thereby deriving our main results. We will prove Theorem 1.2 in Section 3, prove Theorems 1.4 and 1.3 in Section 4, and prove Theorem 1.5 in Section 5.
@@ 2 Notations and Preliminary Results
For a subset $A$ of $V(G)$ and a vertex $b\in V(G)\setminus A$, we say that $b$ is complete to $A$ if $b$ is adjacent to every vertex of $A$, and that $b$ is anticomplete to $A$ if $b$ is not adjacent to any vertex of $A$. For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
@@
3 Proof of Theorem 1.2
Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. We say that a graph $G$ is locally perfect if for every vertex $v\in V(G)$, $G[N(v)]$ is perfect.
In this section, we will prove Theorem 1.2. The following useful lemmas is important to our proof.
@@
[9] Every MNWD graph has no homogeneous set.
@@
[8] If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
@@
Let $G$ be a connected locally perfect bull-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
Proof. Suppose to its contrary that $H$ is an odd antihole of $G[M(v)]$ with $V(H)=\{v_{1},v_{2},...,v_{k}\}$, where $k\geq 7$ and $\overline{H}=v_{1}v_{2}\cdots v_{k}v_{1}$. Let $v^{\prime}\in N(v)$. We first prove that
if $y\in N(V(H))$, then $|N(y)\cap V(H)|\geq 2$. (1)
Suppose not, and we may by symmetry assume that $N(y)\cap V(H)=\{v_{1}\}$. But then $\{v_{1},v_{3},v_{5},y,v_{2}\}$ induces a bull, a contradiction. This proves (1).
We next prove that
if $x\in M(V(H))$ and $y\in N(V(H))$ such that $x\sim y$, then $N(y)\cap V(H)$ is a stable set. (2)
Suppose not, and without loss of generality, we may suppose that $v_{1},v_{n}\in N(y)\cap V(H)$ with $v_{1}v_{n}\in E(G)$, where $3\leq n\leq k-2$. We next show that $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. If not, then let $2\leq n^{\prime}\leq n-1$ be the minimum integer such that $y\not\sim v_{n^{\prime}}$. If $n=3$, then $n^{\prime}=2$. To avoid an induced bull on $\{v_{1},v_{3},y,x,v_{4}\}$, we have that $y\sim v_{4}$; but then $\{v_{1},v_{4},y,x,v_{2}\}$ induces a bull, a contradiction. Hence, $n\geq 4$, and thus $v_{n}\sim v_{2}$ and $v_{n-1}\sim v_{1}$.
We can deduce that $n^{\prime}\neq 2$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{2}\}$; and $n^{\prime}\neq n-1$ to avoid an induced bull on $\{v_{1},v_{n},y,x,v_{n-1}\}$. We have that $3\leq n^{\prime}\leq n-2$, and so $v_{n^{\prime}}\sim v_{n}$ and $v_{1}\sim v_{n^{\prime}-1}$ by the minimality of $n^{\prime}$. But then $\{v_{n^{\prime}-1},v_{n},y,x,v_{n^{\prime}},\}$ induces a bull, a contradiction. Therefore, $y$ is complete to $\{v_{1},v_{2},\cdots,v_{n}\}$. Similarly, with the same arguments, we can deduce that $y$ is complete to $\{v_{1},v_{k},v_{k-1},\cdots,v_{n}\}$, and this implies that $y$ is complete to $V(H)$, which contradicts to that $G$ is locally perfect. This proves (2).
Since $G$ is connected and $V(H)\subseteq M(v)$, there exist two adjacent vertices $x\in M(V(H))$ and $y\in N(V(H))$. By (1) and (2), we may by symmetry assume that $N(y)\cap V(H)=\{v_{1},v_{2}\}$. But then, $\{v_{k-3},v_{k-1},v_{k},v_{1},y\}$ induces a bull, a contradiction. This completes the proof of Lemma 3.3. $\square$
@@
Let $G$ be a bull-free MNWD graph. Then $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect.
Proof. It is certain that $G$ is connected. Since $G$ is MNWD, we have that for every vertex $v\in V(G)$, $G[M(v)]$ cannot be perfect as otherwise, $(\{v\}\cup M(v),N(v))$ is a partition of $V(G)$ such that for all positive integer weight function on $V(G)$, the $G[\{v\}\cup M(v)]$ is a perfect graph and $G[N(v)]$ has a smaller maximum weight of a clique than $G$, which contradicts with the minimality of $G$. Therefore, for every $v\in V(G)$, $G[M(v)]$ is imperfect. Moreover, since $G$ is MNWD, $G$ has no homogeneous set by Lemma 3.1. Therefore, for every $v\in V(G)$, $G[N(v)]$ is perfect by Lemma 3.2. This proves Lemma 3.4. $\square$
By Lemmas 3.3 and 3.4, we have that
@@
Let $G$ be a bull-free MNWD graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
We now prove Theorem 1.2.
Proof of Theorems 1.2: It suffices to show that each (odd torch, bull)-free graph is perfectly weight divisible. Suppose it is not true, and let $G$ be a (odd torch, bull)-free MNWD graph. By Lemma 3.4, we have that $G$ is connected, and
for every vertex $v\in V(G)$, $G[N(v)]$ is perfect and $G[M(v)]$ is imperfect. (3)
Let $v\in V(G)$. By (3), Lemma 3.5 and Theorem 1.1, $G[M(v)]$ contains an odd hole, say $C=v_{1}v_{2}\cdots v_{k}v_{1}$, where $k\geq 5$ is an odd integer. Since $G$ is connected, there exist two adjacent vertices $x\in M(V(C))$ and $y\in N(V(C))$. Since $G$ is odd torch-free, we have that $N(y)\cap V(C)$ is not a stable set. Hence, we may by symmetry assume that $\{v_{1},v_{2}\}\subseteq N(y)\cap V(C)$.
By (3), we have that $y$ is not complete to $V(C)$ as $G[N(y)]$ is perfect. Let $3\leq n\leq k$ be the minimum integer such that $y\not\sim v_{n}$. But then $\{v_{n-2},v_{n-1},y,x,v_{n}\}$ induces a bull, a contradiction. Therefore, each (odd torch, bull)-free graph is perfectly weight divisible. This completes the proof of Theorem 1.2. $\square$
4 Proof of Theorems 1.3 and 1.4
In this section, we will prove Theorems 1.3 and 1.4. The following useful lemmas are important to our proof.
@@
Proof. Let $w$ be a positive integer weight function on $V(G)$, and let $G^{\prime}$ be an induced subgraph of $G$. Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$. Let $X\subseteq B$ such that each vertex in $X$ is a isolated vertex of $G^{\prime}$. Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$. This proves the lemma. $\square$

### v5

# On Perfect Divisibility of Bull-Free Graphs Without Long Paths

@@ Abstract
A graph $G$ is perfectly divisible if, for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. A bull is a graph consisting of a triangle with two disjoint pendant edges. Let $H$ be an odd hole or an odd antihole. An odd net is the graph obtained from $H$ by adding a new vertex adjacent to all vertices of $H$. An odd net+ is the graph obtained from an odd net by adding an additional new vertex that is non-adjacent to any vertex of $H$. Let $F$ denote the Grötzsch graph.
Chudnovsky and Sivaraman [J. Graph Theory 90 (2019) 54-60] proved (bull, $P_{5}$)-free graphs are perfectly divisible. Chen and Xu [Discrete Appl. Math. 372 (2025) 298-307] proved that (bull, $P_{7},C_{5}$)-free graphs are perfectly divisible. In this paper, we prove that (bull, $P_{8},C_{5}$)-free graphs are perfectly divisible. As known $F$ is non-perfectly divisible, and so the class of (bull, $P_{6}$)-free graphs is non-perfectly divisible. Let $G$ be a (bull, $P_{6}$)-free graph. We also show that (i) $G$ is perfectly divisible if and only if $G$ is $F$-free; (ii) if $G$ is odd net-free, then $G$ can be partitioned into two perfect graphs, and so $\chi(G)\leq 2\omega(G)$; (iii) if $G$ is odd net+-free, then $\chi(G)\leq\omega(G)^{2}$; (iv) $\chi(G)\leq\omega(G)^{7}$. These results generalize the findings of Chudnovsky and Sivaraman, and Chen and Xu.
Key words and phrases: bull-free graphs; perfect divisibility; chromatic number.
AMS 2020 Subject Classifications: 05C15, 05C75
@@
Let $H_{1}$ and $H_{2}$ be two vertex disjoint graphs. The join $H_{1}+H_{2}$ is the graph with $V(H_{1}+H_{2})=V(H_{1})\cup V(H_{2})$ and $E(H_{1}+H_{2})=E(H_{1})\cup E(H_{2})\cup\{xy\;|\;x\in V(H_{1}),y\in V(H_{2})$}\}.
For a given positive integer $k$, we use $[k]$ to denote the set $\{1,\ldots,k\}$. A k-coloring of a graph $G=(V,E)$ is a mapping $f$: $V\rightarrow[k]$ such that $f(u)\neq f(v)$ whenever $uv\in E$. We say that $G$ is k-colorable if $G$ admits a $k$-coloring. The chromatic number of $G$, denoted by $\chi(G)$, is the smallest positive integer $k$ such that $G$ is $k$-colorable. A clique (resp. stable set) of $G$ is a set of pairwise adjacent (resp. non-adjacent) vertices in $G$. The clique number of $G$, denoted by $\omega(G)$, is the maximum size of a clique in $G$.
The concept of $\chi$-boundedness was introduced by Gyárfás [15] in 1975. A class of graphs is said to be hereditary if it is closed under isomorphism and induced subgraphs. For a hereditary class $\mathcal{F}$, if there exists a function $f$ such that $\chi(H)\leq f(\omega(H))$ for every $H\in\mathcal{F}$, then $\mathcal{F}$ is called $\chi$-bounded and $f$ a binding function.
An induced cycle of length $k\geq 4$ is called a hole, and $k$ is the length of the hole. A k-hole is a hole of length $k$. A hole is odd if $k$ is odd, and even otherwise. An antihole is the complement graph of a hole. A graph $G$ is said to be perfect if $\chi(H)=\omega(H)$ for every induced subgraph $H$ of $G$. The Strong Perfect Graph Theorem [8] was established by Chudnovsky et al. in 2006:
@@
Let $F$ denote the Grötzsch graph. A bull is a graph consisting of a triangle with two disjoint pendant edges. Let $H$ be an odd hole or an odd antihole. An odd net is the graph obtained from $H$ by adding a new vertex adjacent to all vertices of $H$. An odd net+ is the graph obtained from an odd net by adding an additional new vertex that is non-adjacent to any vertex of $H$. (see Figure 1).
Figure 1: Illustration of some forbidden graphs.
A graph $G$ is perfectly divisible if for each induced subgraph $H$ of $G$, $V(H)$ can be partitioned into $A$ and $B$ such that $H[A]$ is perfect and $\omega(H[B])<\omega(H)$. This concept was proposed by Hoàng in [13]. By a simple induction on $\omega(G)$, we have that every perfectly divisible graph $G$ satisfies $\chi(G)\leq\binom{\omega(G)+1}{2}$ [10].
Let $G$ be a graph and let $w$ be a weight function on $V(G)$. We use $\omega_{w}(G)$ to denote the maximum weight of a clique in $G$. A graph $G$ is perfectly weight divisible [10, 13] if for every positive integer weight function $w$ on $V(G)$, and for every induced subgraph $H$ of $G$, there is a partition of $V(H)$ into $P$ and $W$ such that $H[P]$ is perfect and $\omega_{w}(H[W])<\omega_{w}(H)$. If $G$ is not perfectly weight divisible but each of its proper induced subgraph is, then we say $G$ is minimally non-perfectly weight divisible (MNWD for short). It is certain that if a graph $G$ is perfectly weight divisible, then $G$ is perfectly divisible.
Chudnovsky and Sivaraman [10] proved that every (bull, odd hole)-free graph and every (bull, $P_{5}$)-free graph are perfectly divisible. Karthick et al. [16] proved that every (bull, fork)-free graph is perfectly divisible, where fork is the graph obtained from a $K_{1,3}$ by subdividing an edge once. Chen and Xu [5] proved that every (bull, $P_{7},C_{5}$)-free graph is perfectly divisible. It is certain that $F$ is (bull, $P_{6}$)-free and not perfectly divisible (because $\chi(F)=4$ and $\omega(F)=2$). Actually, the graph $F+K_{n}$ is ($P_{6}$, bull)-free with clique number $n+2$ and not perfectly divisible as $F$ is not perfectly divisible. Further properties of perfectly divisible graphs and bull-free graphs can be found in [6, 12, 14]. Since every (bull, $P_{5}$)-free graph is perfectly divisible, a natural question is to determine the conditions under which a (bull, $P_{6}$)-free graph is perfectly divisible.
In this paper, we generalize the results of Chudnovsky and Sivaraman [10], and Chen and Xu [5] by the following two theorems.
@@
Let $G$ be a $(bull,P_{8},C_{5})$-free graph. Then $G$ is perfectly weight divisible, and hence $G$ is perfectly divisible.
@@
Let $G$ be a $(bull,P_{6})$-free graph. Then $G$ is perfectly divisible if and only if $G$ is $F$-free.
Note that the class of $3K_{1}$-free graphs has no linear binding function [19]. In 2022, Brause et al.[3] gave a more general theorem.
@@
By Theorem 1.4, the classes of $(\text{bull},H)$-free graphs, where $H\in\{\{P_{8},C_{5}\},\{P_{6},F\}\}$, do not have a linear binding function. By Theorems 1.2 and 1.3, we can directly derive the following corollaries by a simple induction on $\omega(G)$.
@@
Let $G$ be an $(\text{bull},H)$-free graph, where $H\in\{\{P_{8},C_{5}\},\{P_{6},F\}\}$ Then $\chi(G)\leq\binom{\omega(G)+1}{2}$.
[…]
@@
Let $G$ be a ((bull, $P_{6}$, odd net))-free graph. Then $G$ can be partitioned into two perfect graphs, and hence $\chi(G)\leq 2\omega(G)$.
Theorem 1.6
[…]
Theorem 1.7
Let $G$ be a $(bull,P_{6})$-free graph. Then $\chi(G)\leq\omega(G)^{7}$.
We will prove Theorem 1.2 in Section 3, prove Theorem 1.3 in Section 4, and prove Theorems 1.5, 1.6 and 1.7 in Section 5.
@@ 2 Notations and Preliminary Results
For a subset $A$ of $V(G)$ and a vertex $u\in V(G)\setminus A$, we say that $u$ is complete to $A$ if $u$ is adjacent to every vertex of $A$, and that $u$ is anticomplete to $A$ if $u$ is not adjacent to any vertex of $A$. For two disjoint subsets $A$ and $B$ of $V(G)$, AA is complete to $B$ if every vertex of $A$ is complete to $B$, and $A$ is anticomplete to $B$ if every vertex of $A$ is anticomplete to $B$.
@@
Let $S\subseteq V(G)$ with $1<|S|<|V(G)|$. We say that $S$ is a homogeneous set of $G$ if every vertex in $V(G)\setminus S$ is either complete to $S$ or anticomplete to $S$. The following useful lemmas are important to our proof.
@@
((Theorem 3.6 of [10])) Every MNWD graph has no homogeneous set.
@@
((4.3 of [9])) If $G$ is a bull-free graph, then either $G$ has a homogeneous set or for every $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
@@
((Lemma 6 of [4])) Let $G$ be a bull-free MNWD graph. Then $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least $7$.
@@
((Theorem 2 of [11])) Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
@@
((Theorem 6 of [18])) Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains $F$ if $\chi(G)=4$.
@@
Proof. Let $w$ be a positive integer weight function on $V(G)$, and let $G^{\prime}$ be an induced subgraph of $G$. Since $\chi(G^{\prime})\leq 3$, there is a partition $(A,B)$ of $V(G^{\prime})$ such that $G^{\prime}[A]$ is bipartite and $B$ is a stable set of $G^{\prime}$. Let $X\subseteq B$ such that each vertex in $X$ is an isolated vertex of $G^{\prime}$. Then $(A\cup X,B\setminus X)$ is a partition of $G^{\prime}$ such that $G[A\cup X]$ is perfect, and $\omega_{w}(G[B\setminus X])<\omega_{w}(G^{\prime})$. This proves the lemma. $\square$
3 Proof of Theorem 1.2
In this section, we prove Theorem 1.2 and begin with the following lemma.